i need to know a step by step waty to get 36 from 3x/4-2=X/2+7
Answer by AnlytcPhil(1810)


i need to know a step by step way to get 36 from 3x/4-2=X/2+7
3x/4 - 2 = x/2 + 7
Multiply every term by LCD of 4
4·(3x/4) - 4·(2) = 4·(x/2) + 4·(7)
3x - 8 = 2x + 28
Add 8 to both sides
3x = 2x + 36
Subtract 2x from both sides
x = 36
Edwin